OpenStudy (anonymous):

The values of x that satisfy the inequality 2x-3/(x+4) <-3 lie in the interval?

OpenStudy (anonymous):

afraid i do not see any inequality here

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

plzz view the question now

OpenStudy (anonymous):

is it \[\frac{2x-3}{x+4}<0\] or maybe \[2x-3<x+4\]?

OpenStudy (anonymous):

\[\frac{2x-3}{x+4}<-3\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first add 3 to both sides to get \[\frac{2x-3}{x+4}+3<0\] then add to get \[\frac{2x-3}{x+4}+3=\frac{2x-3+3(x+4)}{x+4}\] \[=\frac{2x-3+3x+12}{x+4}=\frac{5x-9}{x+4}\]

OpenStudy (anonymous):

ok then...

OpenStudy (anonymous):

now your inequality looks like \[\frac{5x-9}{x+4}<0\] and you should be good to go

OpenStudy (anonymous):

hey but the answer should be (-4 , -9/5)

OpenStudy (anonymous):

find the zeros of each factor. they are -4 and 9/5

OpenStudy (anonymous):

how???

OpenStudy (anonymous):

oh right, because i made a mistake it is 12 -3 =9 not -9 so it should be \[\frac{5x+9}{x+4}<0\] not what i wrote

OpenStudy (anonymous):

ok i want to know how that -4 came........

OpenStudy (anonymous):

that's better. \(x+4=0\) if \(x=-4\) and \(2x+9=0\) if \(x=-\frac{9}{5}\)

OpenStudy (anonymous):

so the quotient \(\frac{5x+9}{x+4}\) will be negative (less than zero) between these numbers, in other words in the interval \((-4,-\frac{9}{5})\)

OpenStudy (anonymous):

so you see where the "-4" came from? it is the value of \(x\) that will make the denominator equal to zero

OpenStudy (anonymous):

if we take x+4 to that side it will be 0

OpenStudy (anonymous):

i am not sure what you mean. if you write \[x+4=0\] you get \[x=-4\]

OpenStudy (anonymous):

and if you set \[5x+9=0\] you get \[x=-\frac{9}{5}\] which i hope explains where the endpoints of the interval come from

OpenStudy (anonymous):

ok i got it......

OpenStudy (anonymous):

you also need to remember that you are asking where this quotient is negative, and it will be negative between the zeros and positive outside them

OpenStudy (anonymous):

ok plzz help @satellite73 new question on maths page http://openstudy.com/study?signup#/updates/4fbf7453e4b0dd370c77304f